OpenStudy (anonymous):

Simplify this Basic Trigonometric Identity: sin A * cot A - Show work -What Identity did you used?

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ cot(\theta)=\cfrac{cos(\theta)}{sin(\theta)}}} \\ \quad \\ \quad \\ sin(A)\cdot cot(A)\implies \cancel{sin(A)}\cdot\cfrac{cos(A)}{\cancel{sin(A)}}\)

OpenStudy (anonymous):

I did not get it! Could you please explain? :(

OpenStudy (anonymous):

@jdoe0001